# Distinguishing Tidal Disruption Events and Changing-look Active Galactic Nuclei via Variation of Mid-infrared Color

### v2

@@ 3.1 Time variation of MIR color in the rising phase
Figure 4: Time variation of MIR color in the rising phase. The upper panel shows TDEs (red) and ANTs (green), and the lower panel shows CLAGNs. For clarity, we split the CLAGN sample into two parts, one with CVR$<0.1$ (blue) and the other with CVR$>0.1$ (dark cyan). We labelled CLAGN Mrk 1018 with special color variation using purple.
@@
As seen from Figure 5, TDEs and ANTs have overall larger CVRs than CLAGNs. The CVRs of 8 TDEs are between 0.27 and 1.11 mag yr-1 with a median of 0.47, and those of 7 ANTs are between 0.20 and 0.56 mag yr-1 with a median of 0.35, and those of 20 CLAGNs are between $-0.55$ and 0.88 mag yr-1 with a median of 0.09. We checked the statistical significance of the difference in CVR distributions across TDEs, ANTs and CLAGNs using the Kolmogorov-Smirnov (K-S) test. The difference between TDEs and CLAGNs (ANTs and CLAGNs) is significant as the probability corresponding to the K-S statistic is $2.9\times 10^{-5}$ ($3.1\times 10^{-4}$). At the same time, there is no significant difference between TDEs and ANTs. In addition, the median IC of 8 ANTs (including AT 2018bcb) is 0.36 mag, smaller than 0.52 mag, the median of 9 TDEs (including AT 2017gge), and the difference between IC distributions is also significant according to the K-S test as the probability is 0.020.
@@ 3.2 Distinguishing TDEs/ANTs from CLAGNs
We first considered the impact of these uncertainties on the maximum of CLAGNs. The sampling leads to underestimation of the maximum, while measurement error leads to overestimation when the sample size is large enough. The two uncertainties lead to opposite effects, whose strengths are related to the sample size, the intrinsic scatter of CLAGNs’ CVRs, and the level of errors. To quantitatively assess which effect is more substantial, we ran a simulation. In the simulation, we assumed that the current observed values were true and evaluated the effects of sampling and measurement error using bootstrap and Monto Carlo methods, respectively. Simulations show that the fake sample overestimated the maximum 97.5% of the time, indicating that the overestimation caused by measurement error dominates the underestimation caused by sampling. According to the simulation, the probability of the fake sample overestimating the maximum or underestimating by $<0.07$ mag yr-1 is as high as 99.3%. Therefore, we considered it safe to set the upper limit of CLAGNs’ CVR at 0.4 mag yr-1.
We then analyzed the minimum CVR of TDE+ANTs similarly. We found that the underestimation of the minimum caused by measurement error dominates the overestimation caused by sampling. Similarly, we adopted a safe lower limit of TDE+ANTs’ CVR at 0.1 mag yr-1, which underestimates the true value by $>99.9\%$ according to the simulations.
@@ 4 Explanations for differences in MIR color variations
In this section, we explore the possible explanations of the differences in MIR color variation between different types of outbursts with the help of simulations. The main phenomena we need to explain are: 1. TDEs’ and ANTs’ MIR colors turn red faster than CLAGNs’ in the rising phase, as the median CVR of TDE+ANTs is $\sim$0.3 mag yr-1 larger than that of CLAGNs; 2. ANTs are bluer than TDEs when they begin, as their median IC is smaller than TDEs.
We simulated the IR LCs of dust echoes of TDEs, ANTs and CLAGNs using a dust radiative transfer model developed by Lu et al. (2016) and Sun et al. (2020). We considered a transient UV source in the galaxy center heats dust grains to generate IR reradiation in a spherical shell. The LC of the transient source is described as $L_{\rm UV}(t)$. The dust shell has an initial inner radius of $r_{\rm in}$ and an outer radius of $r_{\rm out}$, the dust composition is a standard ISM mixture of 47% graphite and 53% silicate, and the dust grain has a radius of $a_{0}$ and a number density of $n_{d}$. The model calculates IR LCs by considering the dust’s inner radius change caused by sublimation, the time delay of IR reradiation due to the light travel time effect, and the absorption of emergent IR emission by the dust structure itself.
We investigated the influence of each parameter on the MIR color variation using the control variable method. We display the influence of different UV LCs in section 4.1, and the influence of different echoing dusts in section 4.2, and finally discuss possible explanations of different CVR and IC in section 4.3.
@@ 4.1.1 Influence of shape of UV LC
$L_{\rm UV}(t)=\begin{cases}L_{\rm max}\frac{t-t_{0}}{\tau_{3}}&\text{ , $t_{0}<t<t_{0}+\tau_{3}$ }\\ L_{\rm max}\left(2-\frac{t-t_{0}}{\tau_{3}}\right)&\text{ , $t_{0}+\tau_{3}<t<t_{0}+2\tau_{3}$ }\\ 0&\text{ , else, }\end{cases}$ (3)
@@
We studied the influence of the shape of UV LC on the MIR color variation when the $L_{\rm max}$ is fixed at $10^{44}$ erg s-1 and $\tau$ is fixed at $10^{7}$ s. In simulations hereafter, we set the default value of initial inner radius $r_{\rm in}$ as $0.6r_{\rm sub}(L_{\rm max})$, where $r_{\rm sub}$ is the sublimation radius corresponding to $L_{\rm max}$, and is calculated as (e.g., Peterson, 1997):
@@
The model can calculate a new inner radius by taking the dust sublimation into account, and hence the value of $r_{\rm in}$ has little influence on the IR LCs when it is less than $0.8r_{\rm sub}$. In addition, the default values of $r_{\rm out}$, $a_{0}$ and $n_{d}$ are 1 pc, 0.1 $\mu$m and $10^{-8}$ cm-3, respectively.
@@
As shown in Figure 6, the MIR color curves in the rising phases are different. For the TDE and EXP forms, initially, the MIR color is relatively blue and becomes red continuously in the rising and fading phases. While for the LINEAR and FLAT forms, the MIR color turns slightly blue or keeps steady with the rapid rise of luminosity in the early time, and then keeps a blue color near the peak, and finally turns red as it fades. We sampled the simulated MIR color curve at 180-day intervals and calculated IC and CVR in the rising phase using the sampled data. IC and CVR are not definite values but in a range because of the uncertain sampling time, expressed as the time delay $t_{\rm delay}$ between the first sampling and the outburst time. We found that the form of LC has little effect on the IC, but affected the CVR: CVR is 0.2–0.4, 0.15–0.35, $-0.1$–0.2 and $\sim 0.05$ mag yr-1 for TDE, EXP, LINEAR and FLAT forms. We also tried other forms of UV LC with more complicated models. By summarizing these results, we found that to produce a large CVR of $\gtrsim 0.3$ mag yr-1, the UV LC needs to contain a rapid rise, a short peak, and a long tail. The reason is that the echoes of the peak luminosity and the long tail have blue and red MIR colors, respectively, and in the rising phase, the fraction of the echo of the long tail is increasing, causing the MIR color to turn red continuously.
@@ 4.1.2 Influence of underlying AGN
ANTs and CLAGNs occur in AGNs with long-lived IR emission from the dusty torus. The underlying AGN causes multiple simultaneous effects, and we explored these effects using simulations. We added a long-standing AGN component with a luminosity $L_{\rm AGN}$ of $10^{43}$$10^{44}$ erg s-1 to the UV LC. We tried UV LC in TDE and LINEAR forms, with $L_{\rm max}$ of $10^{44}$ erg s-1 and $\tau$ of $10^{7}$ s. Note that we subtracted the quiescent fluxes when calculating the MIR color in Section 3, so in the simulation, we also calculated the MIR color after subtraction the IR luminosities of the underlying AGN (the values at $t<=t_{0}$). The results are shown in Figure 7.
@@
In addition, we also explored the influence of the peak UV luminosity and duration on IC and CVR. We show the results in detail in Appendix E.1, and in brief, these two factors could not explain the differences in CVRs and ICs between different types of outbursts.
@@ 4.2.1 Influence by dust properties
The chemical composition and grain radius of the dust influence the MIR color. As shown in Figure B1(a), graphite dust is bluer than silicate dust at the same temperature. Moreover, the sublimation temperature of graphite ($\sim 1900$ K) is higher than that of silicate ($\sim 1500$ K), resulting in a much higher MIR color of graphite ($\sim 0.2$ mag) than silicate ($\sim-0.3$ mag) at around sublimation temperature. Thus, the difference in dust composition can affect the IC: a higher fraction of graphite results in a bluer MIR color (smaller IC). Dust grains with different radii with the same composition show a small difference in color of 0.1–0.2 mag at the same temperature. Thus, the grain radius may slightly affect the IC. Although we have some knowledge of the dust properties in AGNs, those in TDEs are poorly constrained by observations. Future JWST MIR spectroscopic observations may constrain the dust properties in TDEs, check whether they differ from AGNs, and explore if they can explain the observed IC difference between TDEs and ANTs.
@@
Does the difference in dust properties affect CVR? Answering this question requires considering that dust composition may be dependent on locations. Assuming that the initial dust is a mixture of graphite and silicate, the innermost region of the final dust structure would have a high graphite fraction because the sublimation radius of the graphite is larger than that of the silicate. After taking this effect into account, the color difference between the inner region (arrives earlier) and the outer region (arrives later) would be larger, resulting in a larger CVR. However, this effect may work in both TDE and CLAGN, so it might not explain the difference in CVR between the two samples.
4.2.2 Influence by dust shape
The emergent optical depth of the dust structure significantly affects IC according to our simulations (Appendix E.2). The IC increases (redder) with the increase of the emergent optical depth. The difference in the radial optical depth of the dust structure or the difference in the dust geometry may cause a difference in the emergent optical depth. An example of the latter is that the IR echo of a face-on torus, which mainly comes from the inner region, has a smaller emergent optical depth than a spherical shell or edge-on torus with the same radial optical depth.

### v3

@@ 3.1 Time variation of MIR color in the rising phase
Figure 4: Time variation of MIR color in the rising phase. The upper panel shows TDEs (red) and ANTs (green), and the lower panel shows CLAGNs. For clarity, we split the CLAGN sample into two parts, one with CVR<0.1<0.1 (blue) and the other with CVR>0.1>0.1 (dark cyan). We labelled CLAGN Mrk 1018 with special color variation using purple.
@@
As seen from Figure 5, TDEs and ANTs have overall larger CVRs than CLAGNs. The CVRs of 8 TDEs are between 0.27 and 1.11 mag yr-1 with a median of 0.47, and those of 7 ANTs are between 0.20 and 0.56 mag yr-1 with a median of 0.35, and those of 20 CLAGNs are between 0.55-0.55 and 0.88 mag yr-1 with a median of 0.09. We checked the statistical significance of the difference in CVR distributions across TDEs, ANTs and CLAGNs using the K-S test. The difference between TDEs and CLAGNs (ANTs and CLAGNs) is significant as the probability corresponding to the K-S statistic is $2.9\times 10^{-5}$ ($3.1\times 10^{-4}$). At the same time, there is no significant difference between TDEs and ANTs. In addition, the median IC of 8 ANTs (including AT 2018bcb) is 0.36 mag, smaller than 0.52 mag, the median of 9 TDEs (including AT 2017gge), and the difference between IC distributions is also significant according to the K-S test as the probability is 0.020.
@@ 3.2 Distinguishing TDEs/ANTs from CLAGNs
We first considered the impact of these uncertainties on the maximum of CLAGNs. The sampling leads to underestimation of the maximum, while measurement error leads to overestimation when the sample size is large enough. The two uncertainties lead to opposite effects, whose strengths are related to the sample size, the intrinsic scatter of CLAGNs’ CVRs, and the level of errors. To quantitatively assess which effect is more substantial, we ran a simulation. In the simulation, we assumed that the current observed values were true and evaluated the effects of sampling and measurement error using bootstrap and Monto Carlo methods, respectively. Simulations show that the fake sample overestimated the maximum 97.5% of the time, indicating that the overestimation caused by measurement error dominates the underestimation caused by sampling. According to the simulation, the probability of the fake sample overestimating the maximum or underestimating by <0.07<0.07 mag yr-1 is as high as 99.3%. Therefore, we considered it safe to set the upper limit of CLAGNs’ CVR at 0.4 mag yr-1.
We then analyzed the minimum CVR of TDE+ANTs similarly. We found that the underestimation of the minimum caused by measurement error dominates the overestimation caused by sampling. Similarly, we adopted a safe lower limit of TDE+ANTs’ CVR at 0.1 mag yr-1, which underestimates the true value by $99.9\%$ according to the simulations.
@@ 4 Explanations for differences in MIR color variations
In this section, we explore the possible explanations of the differences in MIR color variation between different types of outbursts with the help of simulations. The main phenomena we need to explain are: 1. MIR color of TDEs and ANTs turns red faster than that of CLAGNs in the rising phase, as the CVRs of TDE+ANTs and CLAGNs are $\gtrsim 0.2$ and $\lesssim 0.3$ mag yr-1, respectively; 2. ANTs are bluer than TDEs when they begin, as their median IC is smaller than TDEs.
We simulated the IR LCs of dust echoes of TDEs, ANTs and CLAGNs using a dust radiative transfer model developed by Lu et al. (2016, hereafter Lu16). The model was described in detail in section 3 of Lu16. In brief, it calculates the IR reradiation of a dust spherical shell heated by a transient UV source in the galaxy center. The LC of the transient source is described as $L_{\rm UV}(t)$. The dust shell has an initial inner radius of $r_{\rm in}$ and an outer radius of $r_{\rm out}$. The dust composition is a standard ISM mixture of 47% graphite and 53% silicate. The dust grain has a radius of $a_{0}$, and is uniformly distributed with a number density of $n_{d}$. This model has the advantage in that it calculates the decrease in the dust grain size caused by sublimation by using thermal equilibrium, thereby consistently obtaining how the inner radius of the dust structure increases over time.
For CLAGNs and ANTs, the echoing dust is likely the dusty torus in the AGN unified model (Antonucci, 1993). In order to simulate the echo of torus dust, we developed Lu16’s model as described in the appendix E.1. In the torus dust case, the model has two more free parameters, including the half-opening angle and the inclination angle $i$.
We investigated the influence of each parameter on the MIR color variation using the control variable method. We display the influence of different UV LCs in section 4.1, the influence of different echoing dusts in section 4.2, and finally discuss possible explanations of different CVR and IC in section 4.3.
@@ 4.1.1 Influence of shape of UV LC
[…]
@@
We estimated the typical $L_{\rm max}$ and $\tau$ for TDEs and CLAGNs. The typical $L_{\rm max}$ of TDEs are $10^{43.0}-10^{45.4}$ erg s-1 (e.g., Yao et al., 2023). We calculated the durations $\tau$ of ZTF TDEs to be $\sim 40-200$ days using the LC models of van Velzen et al. (2021). On the other hand, CLAGNs’ $L_{\rm max}$ and $\tau$ range widely. Assuming that $L_{\rm UV}$ is $\sim$10 times $L_{\rm W2}$ according to typical AGNs’ SEDs, the CLAGNs in this work have $L_{\rm max}$ in the range of $\sim 10^{43}-10^{45}$ erg s-1. Most CLAGNs last for years to decades (e.g., Guo et al., 2024), while some last for only a few months, as seen in the 2018 outburst of NGC 1566 (Oknyansky et al., 2019). Thus, in the simulations, we set the default values of $L_{\rm max}=10^{44}$ erg s-1 and $\tau=10^{7}$ s.
[…]
@@
The dust radiative transfer model we used can calculate a new inner radius by taking the dust sublimation into account, and hence the value of $r_{\rm in}$ has little influence on IR LCs. The outer radius $r_{\rm out}$ and grain density $n_{d}$ affect the UV optical depths of the dust shell as:
$\tau_{\rm UV}$ $\approx(r_{\rm out}-r_{\rm sub})\pi a_{0}^{2}n_{d}Q_{\rm UV}$ (7)
[…]
@@
We simulated the MIR LCs and calculated CVRs and ICs for different shapes of UV LC. As shown in Figure 6, the MIR color curves in the rising phases are different. For the TDE and EXP forms, initially, the MIR color is relatively blue and becomes red continuously in the rising and fading phases. While for the LINEAR and FLAT forms, the MIR color turns slightly blue or keeps steady with the rapid rise of luminosity in the early time, and then keeps a blue color near the peak, and finally turns red as it fades. We sampled the simulated MIR color curve at 180-day intervals and calculated IC and CVR in the rising phase using the sampled data. IC and CVR are not definite values but in a range because of the uncertain sampling time, expressed as the time delay $t_{\rm delay}$ between the first sampling and the outburst time. We found that the form of LC has little effect on the IC, but affected the CVR: CVR is 0.2–0.4, 0.15–0.35, 0.1-0.1–0.2 and $\sim 0.05$ mag yr-1 for TDE, EXP, LINEAR and FLAT forms. We also tried other forms of UV LC with more complicated models. By summarizing these results, we found that to produce a large CVR of $\gtrsim 0.3$ mag yr-1, the UV LC needs to contain a rapid rise, a short peak, and a long tail. The reason is that the echoes of the peak luminosity and the long tail have blue and red MIR colors, respectively, and in the rising phase, the fraction of the echo of the long tail is increasing, causing the MIR color to turn red continuously.
@@ 4.1.2 Influence of underlying AGN
ANTs and CLAGNs occur in AGNs with long-lived IR emission from the dusty torus. The underlying AGN causes multiple simultaneous effects, and we explored these effects using simulations. In the simulations, we set the UV light curve of the illuminating source as:
$L_{\rm UV}(t)=L_{\rm flare}(t)+L_{\rm AGN},$ (8)
where $L_{\rm flare}(t)$ is the UV light curve after subtracting the underlying AGN, and $L_{\rm AGN}$ is the UV luminosity of the underlying AGN. For $L_{\rm flare}(t)$, we set TDE and LINEAR form UV LCs for ANTs and CLAGNs, respectively, and set $L_{\rm max}$ of $10^{44}$ erg s-1 and $\tau$ of $10^{7}$ s. We set $L_{\rm AGN}$ to be in the range of $10^{43}$$10^{44}$ erg s-1, corresponding to amplitudes of variability of 0.7–2.6 magnitude, roughly consistent with the observed $\delta W2$ of 0.5–3.2 magnitude. Note that we subtracted the quiescent fluxes when calculating the MIR color in Section 3, so in the simulation, we also calculated the MIR color after subtraction the IR luminosities of the underlying AGN (the values at $t<=t_{0}$). The results are shown in Figure 7, and we also show those with no underlying AGN for a comparison.
@@
4.1.3 Influence of peak UV luminosity and duration
In subsequent simulations, we will focus on two cases: a TDE case with a TDE form $L_{\rm UV}(t)$ and no underlying AGN, and a CLAGN case with a LINEAR form $L_{\rm flare}(t)$ and an underlying AGN with $L_{\rm AGN}=L_{\rm max}$.
Figure 8: Simulations exploring the influence by different $L_{\rm max}$ and $\tau$. Panels (a) and (c) show the MIR color curves (solid and dashed for the TDE and CLAGN cases, respectively), and (b) and (d) show the inferred ICs and CVRs using the sampled data (circles and stars for the TDE and CLAGN cases, respectively). Different colors indicate different $L_{\rm max}$ for panels (a) and (b), and indicate different $\tau$ for panels (c) and (d).
We studied the influence of different UV peak luminosities $L_{\rm max}$ in the range of $10^{43}-10^{45}$ erg s-1 (Figure 8(a), (b)). For both the TDE and CLAGN cases, CVR decreases with the increase of $L_{\rm max}$, while IC is little affected. The $L_{\rm max}$ of TDEs ($10^{43}-10^{45.4}$ erg s-1) are similar to that of CLAGNs in our sample ($\sim 10^{43}-10^{45}$ erg s-1). Thus, the difference in $L_{\rm max}$ can hardly explain that TDEs have higher CVRs than CLAGNs. In addition, the dispersion of $L_{\rm max}$ of about two orders of magnitude may explain the internal CVR dispersion among TDEs, ANTs or CLAGNs.
We then studied the influence of different durations $\tau$ in the range of $10^{6.5}-10^{8}$ s for the TDE case and $10^{7}-10^{8.5}$ s for the CLAGN case (Figure 8(c), (d)). For the TDE case, CVR in the rising phase decreases with the increase of $\tau$. When $\tau$ increases to $\sim 10^{7.5}-10^{8}$ s ($\sim 1-3$ years), the CVR drops to values similar to those in CLAGNs. Fortunately, the durations of most TDEs and ANTs ($\sim 40-200$ days) are not that long, and thus the results obtained by assuming $\tau=10^{7}$ s are not affected. While for the CLAGN form, CVR is little affected by $\tau$.
@@ 4.2.1 Influence by dust properties
The chemical composition and grain size of the dust influence the MIR color. As shown in Figure B1(a), graphite dust is bluer than silicate dust at the same temperature. Moreover, the sublimation temperature of graphite ($\sim 1900$ K) is higher than that of silicate ($\sim 1500$ K), resulting in a much higher MIR color of graphite ($\sim 0.2$ mag) than silicate ($\sim-0.3$ mag) at around sublimation temperature. Thus, the difference in dust composition can affect the IC: a higher fraction of graphite results in a bluer MIR color (smaller IC). Dust grains with different radii with the same composition show a small difference in color of 0.1–0.2 mag at the same temperature. Thus, the grain size may slightly affect the IC.
To explain that ANTs have bluer MIR colors than TDEs, it is required to assume that the echoing dust of ANTs, which is likely the inner region of AGN torus, has a higher graphite fraction and a larger grain size than the echoing dust of TDEs. However, although we have some knowledge of the dust properties in AGNs, those in TDEs are poorly constrained by observations. Future timely follow-up spectroscopies of dusty TDEs may constrain the dust properties, check whether they differ from AGN tori, and explore if they can explain the observed IC difference between TDEs and ANTs.
@@
Does the difference in dust properties affect CVR? Answering this question requires considering that dust composition may be dependent on locations. Assuming that the initial dust is a mixture of graphite and silicate, the innermost region of the final dust structure would have a high graphite fraction because the sublimation radius of the graphite is larger than that of the silicate. This scenario was supported by observations, as Masterson et al. (2025) found that the MIR spectrum of a dusty TDE, WTP14adbjsh, can be explained with a sum of a close graphite dust component and a distant mixed dust component. After taking this effect into account, the color difference between the inner region (arrives earlier) and the outer region (arrives later) would be larger, resulting in a larger CVR. However, this effect may work in both TDE and CLAGN, so it might not explain the difference in CVR between the two samples.
4.2.2 Influence by inner and outer radii
For TDEs and AGNs, the inner boundary of the echoing dust is generally determined by dust sublimation. In this case, it would be better to describe the actual inner radius with $r_{\rm sub}$, which is determined by $L_{\rm max}$ and is not an independent parameter. Thus, the setting of initial inner radius $r_{\rm in}$ in the simulation would have little influence as long as $r_{\rm in}<r_{\rm sub}$. This was confirmed by our simulations where we varied $r_{\rm in}$.
Figure 9: Simulations exploring the influence of the shape of dust structure. Panels (a) and (c) show the MIR color curves (solid and dashed for the TDE and CLAGN cases, respectively), and (b) and (d) show the inferred ICs and CVRs using the sampled data (circles and stars for the TDE and CLAGN cases, respectively). Different colors indicate different $r_{\rm out}$ for panels (a) and (b), and different geometry for panels (c) and (d).
We then studied the influence of the outer radius $r_{\rm out}$ in the range of 0.3–3 pc. We show the simulation results in Figure 9(a), (b). The $r_{\rm out}$ has little effect on CVR, but significantly affects IC. This is because the IR emission in the rising phase mainly comes from the inner region of the dust structure, and the outer region only affects the optical depth of the emergent IR emission. The increase of $r_{\rm out}$ leads to the increase of $\tau_{\rm MIR}$, thus causing the MIR color to turn red. However, this reddening hardly changes over time, and thus has little effect on CVR.
Can the difference in MIR emergent optical depth explain why ANTs have smaller (bluer) ICs than TDEs? As we have demonstrated in section 4.1.1, the echoing dust in TDEs has lower UV optical depth than the AGN dusty torus. However, ANTs have bluer MIR colors than TDEs, suggesting smaller MIR emergent optical depth. These two are contradictory under the assumption of spherical shell. We will consider other dust geometries in the following subsection.
4.2.3 Spherical shell or torus
We studied the influence of CVR and IC by dust geometry (spherical shell or torus) and the inclination angle of the torus. We ran simulations for both types of geometry. For the torus geometry, we considered two inclinations, $30^{\circ}$ for face-on torus and $80^{\circ}$ for edge-on torus. For both cases the half-opening angle was set to be $40^{\circ}$. The results are shown in Figure 9(c), (d).
Face-on torus results in much smaller (bluer) IC than edge-on torus and spherical shell. The reason is that the initial MIR emission, mainly from the innermost dust, emerges with a much smaller optical depth. The difference in IC is positively correlated with the radial optical depth of the spherical shell or edge-on torus. In our simulations where we assumed $\tau_{\rm UV}\sim 20$, the IC difference is approximately 0.25 magnitude.